A 5% solution of a drug is to be mixed with some 15% solution and some 10% solution to get 20 ml of 8% solution. The amount of 5% solution used must be 2 ml more than the sum of the other two solutions. Set up a system of equations that could be used to determine the number of ml of each solution that should be used.
Answer by stanbon(75887)

A 5% solution of a drug is to be mixed with some 15% solution and some 10% solution to get 20 ml of 8% solution. The amount of 5% solution used must be 2 ml more than the sum of the other two solutions. Set up a system of equations that could be used to determine the number of ml of each solution that should be used.
-----------------------
v + f + t = 20
0.05v + 0.15f + 0.10t = 0.08*20
v = f+t+2
----------------------------
Rearrange:
v + f + t = 20
5v + 15f + 10t = 8*20
v - f - t = -2
================================
Cheers,
Stan H.
===========